OpenStudy (anonymous):

write an equation of the line containing the given point and perpendicular to the given line. (0,5);9x+2y=3 the equation of the line is y= ?

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

first find the slope of 9x+2y=3 what do you get?

OpenStudy (anonymous):

-9/2

jigglypuff314 (jigglypuff314):

yep :) then the slope of a line perpendicular to that is the flipped opposite so the flipped opposite would be 2/9 so you need to make an equation of a line that has slope= 2/9 and point (0, 5)

OpenStudy (anonymous):

so how would I do that?

jigglypuff314 (jigglypuff314):

slope-intercept form: y = mx + b when m = slope and b = y-intercept so since slope = m = 2/9 you get y = (2/9)x + b then temporarily plug in point (0, 5) to find " b " so 5 = (2/9)(0) + b solve for b

OpenStudy (anonymous):

5

jigglypuff314 (jigglypuff314):

yup :) so plug back into y = (2/9)x + b so y = (2/9)x + 5

OpenStudy (anonymous):

so do I put the parenthesis?

jigglypuff314 (jigglypuff314):

yeah, it's best if you do since it could be misinterpreted as 2/(9x) otherwise...